OpenStudy (anonymous):

Solve (lnx)^2=3 what is the formula here? D:

OpenStudy (freckles):

can you solve u^2=3 for u?

OpenStudy (anonymous):

√3?

OpenStudy (paxpolaris):

take the sqrt of both sides: \[\ln(x) = \pm \sqrt3\]

OpenStudy (freckles):

so based on what @PaxPolaris says you have two equations to solve

OpenStudy (anonymous):

kay wait so ln(x)=±√3 10^(±√3)=x?

OpenStudy (anonymous):

wait no ln is e ln(x)=±√3 e^(±√3)=x??

OpenStudy (paxpolaris):

correct

OpenStudy (anonymous):

ah thank you

OpenStudy (anonymous):

@PaxPolaris sorry to bother you again but say I have this lnx = lg2 i can't seem to solve it ><

OpenStudy (paxpolaris):

take e^ of both sides

OpenStudy (paxpolaris):

x= e^(lg2)

OpenStudy (anonymous):

what's e^ of both sides? /shot

OpenStudy (anonymous):

like i mean, how do you do that? is it a square or something

OpenStudy (paxpolaris):

\[\huge x=e^{\log \left( 2 \right)}\]

OpenStudy (anonymous):

but what identity/formula are you using?

OpenStudy (paxpolaris):

if ln x= c then x=e^c ... you used this before ... i don't get what you are asking

OpenStudy (anonymous):

oh right haha sorry about that, i couldn't see it for a moment